Calculate NDP at MP, NDP at FC and NNP at


FC


GrossDomestic Product atRs Crores

Market Prices20000

Indirect Taxes2000

Subsidies500

Exports2000

Imports2500

Capital Consumption1000

Ans:

NDP at MP= 19800 Crores

NDP at FC= 17500 Crores

NNP at FC= 17000 Crores


1
Expert's answer
2021-07-06T12:05:50-0400

NDP at MP= Market Prices-Capital Consumption

NDP at MP= 20000 -1000=19000

NDP at MP= 19000 Crores


NDP at FC= NDP at MP - Indirect tax - Subsidies

NDP at FC= 19000-2000+500=17500

NDP at FC= 17500 Crores


NNP at FC= Market Prices - Indirect Taxes+ Subsidies + (Exports - Imports)-Capital Consumption

NNP at FC= 2000-2000+500+(2000-2500)-1000

NNP at FC= 17000 Crores